Plz Explain "MULTINOMIAL THEOREM" with its applications AND a example (if possible) ....

Grade:12

1 Answers

vikas askiitian expert
13 years ago
  • Multinomial theorem means nothing but how to write down a power of a sum in conditions of powers of the terms in that sum.
  • Multinomial theorem is nothing but rule of a sum in term of rules of the addends. 
  • Multinomial theorem is also called a polynomial theorem. This multinomial is the simplification of the binomial theorem to polynomials.

Multinomial Theorem Examples:

                    For example, (a + b + c) 3= a3 + b3 + c3 + 3a2b + 3a2c + 3b2a + 3b2c + 3c2a + 3c2b + 6abc.

                   Now we calculated each coefficient by first expanding (a + b + c) 2 = a2 + b2 + c2 + 2ab + 2bc + 2ac, then self-multiplying it again to get (a + b + c) 3.

                  On the supplementary this is sluggish process, and can be avoided by means of the multinomial theorem. It solves this method by giving us the closed form for any coefficient we might want.

                 It is possible to the multinomial coefficients from the conditions by using the multinomial coefficient formula.

                 Multinomial theorem is nothing but expression sum in term of terms of the addends.  Multinomial theorem is also known as polynomial theorem. This multinomial is the simplification of the binomial theorem to polynomials.

                A multinomial allocation is the option sharing of the outcome from a multinomial experiment. The multinomial method defines the chance of any result from a multinomial experiment.

Examples: 4x3y + z. This is the example of multinomial.